In a competitive market, the market demand is Qd= 60 −6Pand the market supply is Qs= 4P. A price ceiling of $4 will result in  

(a) A shortage of 18 units.

  1. A shortage of 20 units.
  2. A surplus of 30 units.
  3. A surplus of 12 units.         
  4. neither a shortage nor a surplus.
Answer #1

Answer is option B)

At P = 4,

Qd = 60-6*4

= 60-24

= 36

& Qs= 4*4 = 16

Thus as Qd > Qs

Thus shortage exists , hence c & d, e are false.

now shortage, = Qd- Qs

= 36-16

= 20 units Shortage